OpenStudy (anonymous):

Assuming x ≠ 0 and y ≠ 0, what is the quotient of

OpenStudy (anonymous):

\[\frac{ 12x ^{6} y ^{2}+8x ^{4}y ^{3}+4x ^{2}y ^{4}}{ 4x ^{2}y }\]

ganeshie8 (ganeshie8):

i think you can factor 4x^2y in numerator, then the denominator cancels compeltely... try..

OpenStudy (anonymous):

um how do you factor again?

ganeshie8 (ganeshie8):

numerator : 4 is common in all terms, right ?

OpenStudy (anonymous):

yes ithink so

ganeshie8 (ganeshie8):

\(12x^6y^2+8x^4y^3+4x^2y^4\)

ganeshie8 (ganeshie8):

= \(4(3x^6y^2+2x^4y^3+x^2y^4)\)

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

yes? :)

ganeshie8 (ganeshie8):

\(4y(x^6y+8x^4y^2+4x^2y^3)\)

ganeshie8 (ganeshie8):

\(4x^2y(x^4y+8x^2y^2+4y^3)\)

ganeshie8 (ganeshie8):

does that make sense how we factored 4x^2y ?

OpenStudy (anonymous):

i think so..:P yes

ganeshie8 (ganeshie8):

great ! so, \(\huge \frac{ 12x ^{6} y ^{2}+8x ^{4}y ^{3}+4x ^{2}y ^{4}}{ 4x ^{2}y } \) = \(\huge \frac{4x^2y(3x^4y + 2x^2y^2 + y^3)}{4x^2y} \) = \(\huge \frac{\cancel{4x^2y}(3x^4y + 2x^2y^2 + y^3)}{\cancel{4x^2y}} \) = \(\huge 3x^4y + 2x^2y^2 + y^3\)

ganeshie8 (ganeshie8):

does that help/make sense @luvbouncer11

OpenStudy (anonymous):

yes :) thanks